OpenStudy (sophadof):

Given the ph of 7 explain how many hydroxide -OH and hydronium H30+ are present in the solution

OpenStudy (anonymous):

It depends on the temperature, because Kw can vary. At SATP their concentrations are both equal to 1.0 x 10^{-7}.

OpenStudy (sophadof):

It didn't say anything else but that...

OpenStudy (anonymous):

Then go with 1.0 x 10^{-7} for each.

OpenStudy (sophadof):

but I don't know how to do this stuff...at all :(

OpenStudy (australopithecus):

So, \[pH = -\log_{10}[H+]\] \[[H3O+] = 10^{-pH}\] This gives you the concentration of Hydroxide and Hydronium ions present in a solution at 7pH

OpenStudy (australopithecus):

Well from both of these equations you can find Hydronium (H3O+) and hydroxide (OH) concentrations

OpenStudy (australopithecus):

Oh and the concentration is in Molarity (mol/L) hence the square brakets

OpenStudy (australopithecus):

you can convert mols to particles by using avagadros number which will give you particles per liter and an answer to your question (if you were given a volume you can figure out by multiplication how many particles of each you have in it)

OpenStudy (sophadof):

oh okay!!!(:

OpenStudy (australopithecus):

OOPs

OpenStudy (australopithecus):

Use only the bottom equation I gave you

OpenStudy (sophadof):

the H3O one?

OpenStudy (australopithecus):

You still need to find -OH in solution which is easy by the following relations pH + pOH = 14 so you can find pOH then use pOH to find molarity of -OH with the following formula \[pOH = -\log_{10}[-OH]\]

OpenStudy (australopithecus):

yeah the H3O one

OpenStudy (australopithecus):

After you find -OH concentration just do the same thing I described above to find particles of -OH

OpenStudy (sophadof):

ahh! okay!! thank you!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!(:

OpenStudy (australopithecus):

No problem if you get lost let me know, getting off my bus now so I wont be around for a few hours, but I will check back later so post any problems if you run into them. You need to use algebra to solve for the concentrations but I have faith you can here have the logarithmic rules for help, http://www.mathwords.com/l/logarithm_rules.htm

OpenStudy (sophadof):

@Australopithecus Im so confused :( I don't have it

OpenStudy (australopithecus):

[H3O+]=10^(−pH) becomes [H3O+] = 10^(-7)M =10^(-7)mol/L pH + pOH = 14 7 + pOH = 14 pOH = 7 pOH=−log10[−OH] [−OH] = 10^(-pOH) [-OH] = 10^(-7)M = 10^(-7)mol/L Avogadro number is 6.022141*10^(23)particles/mol you can find particles/L by multiplication (units cancel) if you have a volume for the solution you can figure out how many OH and H3O particles are present if you multiply particles/L by the volume of the solution Number of -OH equals number of H3O also you should notice so yeah